OpenStudy (anonymous):

Solve the proportion 2 times x all over 5 times x plus 6 equals 2 over x.. What is the value of x? @Nurali

OpenStudy (anonymous):

@ash2326

OpenStudy (ash2326):

We have \[\frac{2x}{5x+6}=\frac{2}{x}\] Is this your question @unheard

OpenStudy (goformit100):

Take cross multiplication in use

OpenStudy (goformit100):

 Solve using Quadratic Equation.